In an experiment, the following readings were observed;

Volume of alcohol flowing per minute= 10^-5m^3

Density of alcohol = 800kg/m^3

Length of tube was 0.5m

Radius of tube is 0.05cm and the height of alcohol is 60cm. Calculate the coefficient of viscosity of alc ohol


1
Expert's answer
2021-07-13T13:34:00-0400

According to Poiseouille equation, the volumetric flow rate of the fluid through long cylindrical pipe of constant cross section is "Q = \frac{\pi R^4}{8 \eta} \frac{\Delta p}{L}", where "L" is the length of the pipe, "R" - radius of the pipe, "\eta" - viscosity of the fluid and "\Delta p" is the pressure drop. Here, the pressure drop is "\Delta p = \rho g L", where "\rho" is the density of the fluid.

Combining last two formulas, obtain:

"Q = \frac{\pi R^4}{8 \eta} \frac{\rho g L}{L} = \frac{\pi R^4}{8 \eta} \rho g", from where the viscosity of the alcohol is "\eta = \frac{\pi R^4 \rho g}{8 Q} \approx 0.193 Pa \cdot s".